Fermi National Laboratory

Accelerator Update for November 5, 2008 - November 7, 2008


Wednesday Nov 5

The day shift began with the Operations monitoring the Tevatron (TeV) store 6541 (68.74E30), the Recycler stash (177.3E10), and the Antiproton Source (Pbar) stack (23.1E10), with Linac having RF problems, with Booster in access to replace an RF power amplifier, with the Main Injector (MI) operating normally, with the MiniBooNE and NuMI experiments waiting for the Linac problems to get resolved, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

Operations reported at 8:16 AM that the breaker for a Linac RF station (LRF4) had been replaced.

Booster came out of access at 9:53 AM.

At 10:22 AM, the LRF4 breaker tripped off again. Operations contacted the Duty Mechanic.

A Pbar expert began a cooling study at 11:12 AM.

The Neutron Therapy Facility reported at 12:52 PM that they had treated their last patient for the day.

At 2:34 PM, Fermilab suffered a power outage. Nearly everything but MI and the Recycler had tripped off. Later, it was reported that a static line had torn lose and shorted one of the main 345 thousand volt (kV) power lines coming into the Master Substation. The circled item in the picture below showed the bracket that should have been holding the static line to the top of left pole.

High voltage experts had power back to site through the Kautz Road Substation at 3:37 PM.

At 4:17 PM, Operations reported that operators had begun restoring CAMAC crates.

Operations prepared the TeV for access at 4:27 PM. Other areas being prepared for access included Linac and MI.

Operations reported at 5:11 PM that CUB was fully operational. Operators turned on the TeV LCW system.

Operations prepared D0 and CDF for access around 5:43 PM.

A controls expert reported at 5:52 PM that the Meson Front End (FE) was up and running.

Operations began the TeV recovery at 6:57 PM.

At 8:41 PM, Operations reported a TeV sector B49 spool piece LCW leak.

Operations received word at 10:36 PM that the static line had been repaired and remounted. All other damage was repaired in place. High voltage experts began the static line mount inspections.

Thursday Nov 6

The midnight shift began with Fermi personnel, the City of Batavia personnel, and out side contractors helping to recovering from the 345 kV feeder fault and broken static line.

At 12:51 AM, Operations reported that the beam tube vacuum was bad at TeV sectors A2, B4, C1, and F4. Also, the cryo system vacuum was bad at A4, B1, B2, B3, C3, D4, and F2. Vacuum experts had been adding extra vacuum pumps to all of those areas.

Operations reported at 1:42 AM that high voltage contractors had inspected all the static line mounts on all the poles. They discovered one bracket with a missing cotter key and three others with loose keys. They tightened everything as they conducted their inspection.

A Fermi high voltage expert reported at 2:18 AM that all 345 kV static line work and inspections were complete.

Operations prepare Pbar for an access at 10:05 AM.

Operations prepared MiniBooNE and Booster for access at 10:44 AM.

LCW experts reported at 5:23 PM that they had repaired all Pbar transport LCW leaks.

Operators reported at 5:29 PM that they had search and secured the Switchyard (SY) enclosures.

The TeV came out of access at 5:44 PM.

Operators put the TeV separators into conditioning mode at 6:04 PM.

Pbar came out of access at 7:21 PM.

Friday Nov 7

The midnight shift began with the Fermi accelerator complex in shutdown.

Operators brought the TeV separators out of condition mode at 4:33 AM.

At 5 AM, a planned power outage to restore power through the Master Substation began. The power outage lasted 12 minutes.

Operators began restoring CAMAC Front Ends at 5:21 AM.

Operators restarted all the TeV LCW pumps at 5:33 AM.

Operators reported at 7:03 AM that they were turning on Linac and Booster.

The Plans for Friday and the Weekend

The plans for today and the weekend are to bring the accelerator complex and its systems back online. The Run Coordinator expects beam to be established to SY, NuMI, and MiniBooNE by this afternoon. The TeV should be cold by Sunday. Hopefully, the next store will go into the TeV on Monday, 11/10/08.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 11/07/2008    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory